OpenStudy (anonymous):

HELP! MEDAL WILL BE GIVEN! A material decays at a rate of 1.1% per year. If you start with 250 grams of the material, how much will be left in 15 years? Round to the nearest whole gram. Write a equation that correctly models the situation!

OpenStudy (anonymous):

@HELP!!!!

OpenStudy (anonymous):

what are the choices?

OpenStudy (mathmate):

It's like the compound interest formula, but with a negative rate (of interest).

OpenStudy (anonymous):

There are no choices:( But it should look something like this: A=250(.011)^15 <-- But that is not correct.

OpenStudy (jessiegonzales):

A{n} = 250(1-.011)^(15) = 211.7797

OpenStudy (mathmate):

Compound interest formula is \(A=P(1+i)^t\)

OpenStudy (anonymous):

@jessiegonzales is correct

OpenStudy (anonymous):

@jessiegonzales That looks right! Thank you so much!

OpenStudy (jessiegonzales):

YW :)